Calculate the moles of potassium sulfide formed when 5.4 g of potassium gas reacts with excess sulfur.


2 K+S > K2S


1
Expert's answer
2021-05-08T23:33:56-0400

M(K) = 39.1 g/mol

"n(K) = \frac{m}{M} = \frac{5.4}{39.1}= 0.138 \;mol"

According to the reaction:

n(K2S) "= \frac{1}{2}n(K) = \frac{1}{2} \times 0.138 = 0.069 \;mol"

Answer: 0.069 mol